### v3

# Electron-phonon-dominated charge-density-wave fluctuations in TiSe2 accessed by ultrafast nonequilibrium dynamics

Abstract
The complex phase diagram of 1T-TiSe2 consists of a charge density wave (CDW) below 200 K, and CDW fluctuations of still unknown origin at higher temperatures. Here, we use time-resolved extreme ultraviolet momentum microscopy and density functional perturbation theory to uncover the formation mechanism of CDW fluctuations and their spectral features at 295 K. We investigated the transient dynamics of fluctuations upon nonresonant ultrafast photoexcitation, and directly correlate it with the CDW soft-phonon hardening. Surprisingly, our results show that the coherent amplitude mode modulating ultrafast CDW recovery persists above , and reveal that CDW fluctuations are dominated by the electron-phonon interaction rather than excitonic correlations as commonly believed. Our findings on these microscopic CDW fluctuations clarify the complex interplay between electronic and lattice degrees of freedom at elevated temperatures and, therefore, could be useful in understanding the nature of the CDW phase transition in 1T-TiSe2 and similar quantum materials.
Introduction
Understanding the interplay between electronic and lattice degrees of freedom is essential for comprehending phase transitions, superconductivity, transport properties, and the underlying mechanism of complex quantum phases. Disentangling the origin of the charge density wave (CDW) phase in 1T-TiSe2, with K and a lattice reconstruction Di Salvo et al. (1976); Woo et al. (1976a), provides the right platform to do so. Despite extensive experimental and theoretical investigations, the driving mechanism of the CDW phase remains a debated question Rossnagel (2011). One perspective attributes the origin to electron-hole coupling Jérome et al. (1967), as in the excitonic insulator scenario Di Salvo et al. (1976); Kidd et al. (2002); Cercellier et al. (2007); Monney et al. (2009, 2010, 2012); Kogar et al. (2017), while alternative argues that the dominant mechanism is electron-phonon coupling, represented by the band-type Jahn-Teller effect Hughes (1977); Yoshida and Motizuki (1980); Whangbo and Canadell (1992); Rossnagel et al. (2002); Weber et al. (2011). Recently, a series of ultrafast time-resolved studies made an effort to disentangle the two contributions Porer et al. (2014); Monney et al. (2016); Karam et al. (2018); Hedayat et al. (2019); Lian et al. (2020); Otto et al. (2021); Cheng et al. (2022); Heinrich et al. (2023), and reach a consensus that electron-hole and electron-lattice interactions are intertwined in driving the CDW transition van Wezel et al. (2010). The ultrafast femtosecond dynamics (i.e., melting and recovery) of the order parameter (e.g., the CDW gap) Rohwer et al. (2011); Hellmann et al. (2012) was used as an argument for the former, while the laser-induced coherent phonon motion of the CDW amplitude mode Porer et al. (2014); Monney et al. (2016); Hedayat et al. (2019); Duan et al. (2021) for the latter interaction.
Several studies have reported that the CDW phase persists above due to the fluctuation of the order parameter related to enigmatic CDW fluctuations Woo et al. (1976b); Miyahara et al. (1995); Holt et al. (2001); Kidd et al. (2002); Chen et al. (2016); Cercellier et al. (2007); Monney et al. (2010, 2016), where the presence of finite correlation lengths leads to phase incoherent CDW domains across the sample. Despite recent reports discussing the CDW fluctuations in 1T-TiSe2 at such elevated temperatures Mizukoshi et al. (2023); Zhang et al. (2022); Cheng et al. (2022); Guo et al. (2025) with some reporting CDW fluctuation signatures in static angle-resolved photoemission spectroscopy (ARPES) Jaouen et al. (2019), none of these directly mapped the nonequilibrium light-induced dynamics of these fluctuations in energy and momentum space. Applying the latter technique to explore this exotic phase could provide fertile ground to scrutinize the interplay and robustness of electron-hole and electron-phonon interactions at elevated temperatures. Indeed, a study using time-resolved ARPES (tr-ARPES) demonstrated sub-100 fs melting of the fluctuating phase, which was used as an evidence for the exciton-driven CDW fluctuations Monney et al. (2016). Contrary to these claims, a recent time-dependent density functional theory study showed that exciton order is melted above Lian et al. (2019). Moreover, the coherent oscillation of the CDW amplitude mode was observed under ambient conditions using time-resolved optical spectroscopy Mizukoshi et al. (2023), suggesting a lattice involvement in the dynamics of CDW fluctuations. As evidenced by these opposing claims, the CDW fluctuating phase in 1T-TiSe2 is still an elusive puzzle, while it would be highly valuable to understand its origin considering its recently discovered connection to unconvetional superconductivity Joe et al. (2014); Jaouen et al. (2019); Hinlopen et al. (2024), as well as its resemblance to the fluctuations preceding ordered phases (e.g., pseudogap) in other transition metal dichalcogenides (TMDCs) Borisenko et al. (2008) and strongly correlated materials, such as cuprates Arpaia et al. (2019) and kagome metals Zhong et al. (2024); Liu et al. (2025). Unraveling nonequilibrium properties of CDW fluctuations in TiSe2 would also be highly useful for comprehending recently discovered light-induced hidden phases in TiSe2 Duan et al. (2023); Huber et al. (2024).
In this work, we resolve the debate on the origin of CDW fluctuations in 1T-TiSe2 by using time-resolved extreme ultraviolet (XUV) momentum microscopy, which allows ultrafast light-induced dynamics directly in energy-momentum space. The band structure mapping reveals that the backfolded Se 4p band associated with the CDW phase remains detectable even at room temperature (RT), in agreement with previous reports Kidd et al. (2002); Cercellier et al. (2007); Jaouen et al. (2019). We also measured the partial melting of these CDW fluctuations on an ultrafast timescale, followed by a rapid recovery ( fs). This ultrafast melting signal of CDW fluctuations is offset from the maximum of the hot electron temperature by around fs, and overlaps with the population dynamics of excited Ti 3d states. Together, this demonstrates that CDW fluctuations are able to withstand a strong laser excitation and are not directly linked to the hot electron temperature, suggesting an involvement of the phonon mechanism. Interestingly, we also observe that the recovery of the CDW fluctuations is modulated by coherent CDW amplitude phonon modes. We challenge the belief that the fluctuating phase is ruled solely by the excitonic correlations Monney et al. (2016) by reporting the periodic modulations of the backfolded band spectral weight at RT together with density-functional-theory (DFT) based calculations, indicating not only the preserved electron-phonon coupling mechanism at elevated temperatures but also confirming this to be the main contributor in the hybrid mechanism. More specifically, our calculations incorporating the dynamical electron-phonon scattering confirm that the strongly-coupled CDW-related soft phonon is relevant for the observed equilibrium and transient spectral features of CDW fluctuations. Namely, strong interaction between the soft (yet stable) phonon and Se 4p states leads to the fluctuating replica, while light-induced excitations harden the phonon and screen this interaction, consequently quenching the electron-phonon-driven CDW fluctuations. More generally, our study poses a series of strong arguments favoring the electron-phonon mechanism of CDW fluctuation formation at elevated temperature, which also puts a strong emphasis on the electron-phonon scenario in the debate on the origin of the regular CDW phase. Besides the fundamental importance, correct interpretation of this mechanism could unleash the full potential of unconventional CDW-related features, such as CDW chirality Ishioka et al. (2010), tunable nonlinear optical response Zhang et al. (2022), optical gyrotropy Xu et al. (2020), and superconductivity Morosan et al. (2006)
Results
Equilibrium and Nonequilibrium Electronic Band Structure Mapping. The experiments were performed at the Centre Lasers Intenses et Applications (CELIA) in Bordeaux, France. The time-resolved momentum microscopy setup featuring a polarization-tunable infrared (1.2 eV, sub-150) pump and XUV probe (21.6 eV, sub-50) pulses [Fig. 1(a)] allows for mapping of the occupied and unoccupied states (up to 1.2 eV above the Fermi level) over the full photoemission horizon [Fig. 1(b)]. More details about the experimental setup can be found elsewhere Fragkos et al. (2025); Tkach and Schönhense (2025); Tkach et al. (2024a) and in the Methods section.
To perform photoexcited band mapping, we sit at the pump-probe temporal overlap (s-polarized pump pulse) and measure the photoemission intensity while continuously modulating the XUV linear polarization axis angle (going from s- to p-polarization through all intermediate linear polarization axis angles). The three-dimensional photoemission intensity shown in Fig. 1(b) is obtained by summing over all XUV polarization states, allowing for minimized spectral weight suppression caused by photoemission matrix-element effects Moser (2017); Schusser et al. (2024, 2022); Fragkos et al. (2025). The low-energy electronic band structure of 1T-TiSe2 is characterized by a Se 4p valence band located at the (A) point and Ti 3d conduction band pocket at the M(L) point. As shown in Fig. 1(b), one clear indication of the CDW phase in static ARPES is the backfolded Se 4p band at the M point Kidd et al. (2002); Cercellier et al. (2007). Figure 1(c) shows a comparison between experimental and theoretical constant energy contours (CEC), near the Fermi level. The measured loss of spectral weight within M pockets (black arrows) at RT is another fingerprint of the CDW fluctuations, and it is well captured by theoretical calculations, which include dynamical electron-phonon interaction for temperatures above . Detailed analysis of the theoretical calculations will be given later. These scattering hot spots are reminiscent of the pseudogap openings and loss of spectral weight at and certain momenta relevant for the CDW in other TMDCs, such as 2H-TaSe2 Borisenko et al. (2008) and 2H-NbSe2 Borisenko et al. (2009).
To gain further insight into the signatures of RT-CDW fluctuations, we perform an in-depth analysis of the spectral features around the M(L) pocket for unexcited [Figs. 2 (a)-(b)] and photoexcited 1T-TiSe2 (at pump-probe temporal overlap) [Figs. 2 (c)-(d)]. The photoemission intensity along the (A) - M(L) cut [Fig. 2(a)] reveal spectral features associated with RT-CDW fluctuations. First, we note the presence of the well-established backfolded Se 4p valence band, typically used as the ARPES evidence of the emergence of CDW order, exhibited by a faint yet discernible down-dispersive band feature at the M(L) point below the Ti 3d band in Fig. 2(a). This feature is more apparent in the momentum distribution curves (MDC) around Å-1, as indicated by the black bars in Fig. 2(b). The distance between the two peaks increases with higher binding energies, demonstrating the hole-like Se 4p band character. The backfolded Se 4p feature is also observed at pump-probe overlap in both the photoemission intensity image and MDC cut [Figs. 2(c) and (d)], indicating that CDW fluctuations are not completely melted and persist upon photoexcitation at this relatively high fluence (0.95 mJ/cm2). Additionally, a first-principles theoretical study of the electronic response demonstrated that the excitonic mode is quenched upon heating above Lian et al. (2019), while ultrafast pump-probe optical spectroscopy and tr-ARPES studies of 1T-TiSe2 in the CDW regime () argued that the excitonic correlations should be destroyed above the critical fluence of 0.04 mJ/cm2 Porer et al. (2014); Huber et al. (2022); Cheng et al. (2022). This robustness of the fluctuating replica band at RT and high fluence conditions suggests that the CDW fluctuations are not solely of excitonic origin.
Ultrafast Melting and Recovery of RT-CDW Fluctuations. Now that we have established the equilibrium spectral signatures associated with RT-CDW fluctuations, we turn our attention to their ultrafast melting and recovery upon photoexcitation to gain detailed insight into the microscopic origin underlying the formation of this exotic phase Boschini et al. (2024). The tr-ARPES enables systematic temporal separation of fundamental electronic and structural processes, allowing the identification of the dominant interactions Hellmann et al. (2012).
In this spirit, we record the three-dimensional photoemission intensity as a function of pump-probe delay (between -0.5 ps to 3 ps), yielding a four-dimensional dataset . From this multidimensional dataset, one can hunt for subtle signatures of RT-CDW fluctuations melting and recovery along energy, momentum, and time axes. We start by looking at the differential (pre-time-zero signal subtracted) CEC near the Fermi level ( = -0.07 eV), at a time delay of [Figs. 3 (a) and (b)]. These CECs reveal the population of the conduction Ti 3d band (hollow elliptical positive differential signal) and suppression of the CDW backfolded band spectral weight (central negative differential signal) upon photoexcitation, around M(L) pockets. At this binding energy, after 400 fs, the negative differential signal associated with the melting of CDW order vanishes, as the filling of the bottom of the conduction band becomes dominant. This behavior is also evidenced in the differential energy-momentum region around the L(M) point, extracted for the same pump-probe delays.
To get more insight into the concomitant dynamics of hot electrons and CDW melting/recovery, we analyze the time-resolved photoemission intensities [Fig. 3(c)] in selected regions-of-interest (ROIs), shown as colored boxes in Fig. 3(b). The purple dots represent the IR/XUV cross-correlation extracted from the time-dependent laser-assisted photoemission signal (ROI not shown, cross-correlation shown in Supplementary Figure 6), with an FWHM duration of 171 fs (Gaussian fit, yellow line). The dark red curve represents hot electron populations in Ti 3d conduction band, showing a fast light-induced rise followed by a single exponential decay behavior. The timescale of this relaxation process is strongly dependent on the excess energy of these hot electrons, as shown in the Supplementary Figure 1 in Supplementary Information (SI). As discussed in the context of the analysis of the CECs, the black ROI is spectrally congested, with some overlapping contribution of both the bottom of the Ti 3d conduction band and the top of the Se 4p CDW backfolded band. The dynamics in this ROI feature a fast decay (experimental-resolution-limited) associated with the melting of CDW fluctuations, followed by a refilling and subsequent decay associated with the intraband hot electron dynamics.
To track the CDW fluctuation melting and recovery dynamics in a background-free manner, i.e., in an energy-momentum area where there is no overlap with the spectral feature from the normal phase, we look at the time-resolved photoemission intensity in the blue ROI, which emerges solely due to CDW-fluctuation-driven backfolded bands. The dynamics in this ROI is characterized by an ultrafast melting on a timescale limited by our temporal resolution (170 fs), followed by a recovery on a sub-picosecond timescale, i.e., fs. On top of the sub-picosecond recovery dynamics, some (low contrast) periodic modulations of the photoemission intensity can be observed with a long lifetime. To isolate the contribution from these periodic oscillations, we subtracted the single exponential fitted function from the experimental signal, yielding the residual intensity shown in Fig. 3(d). By performing Fourier analysis, we found that the CDW fluctuations recovery dynamics is periodically modulated by a mode around 3.5 THz. This mode was previously observed in time-resolved photoemission measurements in 1T-TiSe2 below Porer et al. (2014); Monney et al. (2016); Hedayat et al. (2019); Duan et al. (2021) and has been identified as the CDW amplitude phonon mode corresponding to displacements along the CDW periodic lattice distortions. This represents the first time-resolved photoemission observation of coherent amplitude phonon mode surviving in the CDW-fluctuations regime at a temperature strongly exceeding the transition temperature and characterized by an unusually long lifetime.
Another interesting feature in the time-resolved photoemission signal is evident in panel 3(e), where we compare the dynamics of the backfolded Se 4p band with the time evolution of the electronic temperature (extracted from the Fermi-Dirac distribution fitting) and of the total excited state population taken from the yellow ROI in Fig. 3(b), representing the excitation and thermalization of Ti 3d states in a wider energy and momentum region. Note that the increase in electronic temperature and excited-state population observed at negative time delays is attributed to a weak prepulse extending up to a few hundred femtoseconds in the temporal profile of the infrared pump pulses. More importantly, the minimum of the backfolded band intensity is offset from the maximum of the electronic temperature, while it approximately follows the early dynamics of the excited state population of Ti 3d bands. This comparison suggests a mechanism beyond the excitonic insulator scenario, considering that the excitonic order is expected to be melted when the electron temperature reaches its maximum. Namely, exciton, if it exists, would be formed from the holes in Se 4p and electrons in Ti 3d bands, and, therefore, its quenching should be correlated with the total photo-excited carrier density, which can be described with the effective temperature. Furthermore, it is expected that the suppression of the excitonic mode is very fast and that it should correspond to the fastest signal, however, the RT-CDW signal suppression is offset from by approximately fs.
The concomitant observation of partial melting under relatively high-fluence condition (0.95 mJ/cm2), the finite delay between the peak of transient electronic temperature and quenching of the backfolded Se 4p band, as well as coherent CDW amplitude phonon mode modulating the recovery dynamics, supports the idea of the dominant phonon contribution in the assumed hybrid exciton and electron-phonon mechanisms at such elevated temperatures. The idea of dominant phonon contribution is further supported by a recent study Buchberger et al. (2025) which shows that while the band gap is sensitive to substrate screening, the temperature-dependent evolution of its electronic structure remains unaffected, indicating that excitons are not a primary mechanism that drives the CDW transition in TiSe2.
Discussion
To further elaborate on the physical mechanism of CDW-fluctuation melting, two possible suppression pathways can be considered: (i) incoherent suppression, in which coherence is abruptly destroyed by hot electrons on a fast timescale (– fs), and (ii) coherent suppression, mediated by the excitation of the amplitude mode. As previously noted, the comparison of timescales for the maximum suppression of the backfolded band in Fig. 3(e) reveals a near coincidence with the peak in the excited-state population above the Fermi level. Both are delayed relative to the peak in electronic temperature by approximately fs, which closely corresponds to half the period of the amplitude mode oscillation. While it is likely that both coherent and incoherent mechanisms contribute to the suppression, this characteristic time delay—matching half an amplitude mode cycle—indicates a dominant role for coherent suppression.
Further insight can be gained by comparing the time-resolved dynamics of the backfolded band with the timescale of signal suppression at the M point in RT ultrafast electron diffraction Otto et al. (2021). It was shown that the diffuse signal at M point is quenched within the first 130 fs (limited by the instrument response function), and subsequently recovered within the time scale of 700 fs, which remarkably follows the melting time scales of backfolded Se 4p band observed here. The latter quench of the M point diffuse intensity was attributed to hardening of the M acoustic phonon mode by the photo-excited carriers, following the same time dynamics. This directly relates the quenching of the CDW fluctuating band with the hardening of the CDW-related M acoustic phonon.
Another recent ultrafast low-energy electron diffuse scattering study investigated the melting process of charge order in TiSe2 and estimated roughly 30:70 ratio of excitonic versus electron-phonon contributions to the total lattice distortion in the low-temperature CDW phase Kurtz et al. (2024). These findings highlight that both excitonic effects and electron-phonon coupling play significant roles in stabilizing the CDW order. While short-range CDW fluctuations are a dynamical phenomenon—fluctuating over both time and space—the mechanisms that govern their stabilization and ultrafast dynamics are similar to those in conventional low-temperature CDW phases. These short-range CDW fluctuations and their coexistence with some normal (non-CDW) domains can explain the low contrast of the coherent oscillations associated with the CDW amplitude phonon mode. Indeed, in low-temperature standard CDW phases, high-contrast modulations were measured by tr-ARPES Monney et al. (2016); Hedayat et al. (2019); Duan et al. (2021).
Theoretical Analysis of Dynamical Electron-Phonon Interactions. To further disentangle the influence of the electron-phonon coupling on the bare electronic structure above the transition temperature, we turn our attention to the theoretical framework that includes electron self-energy corrections necessary for the inclusion of electron-phonon interaction. Notably, no term responsible for the electron-hole coupling was used here. Yet, the spectral function strikingly matches the experimentally measured features that we attribute to the CDW-fluctuating phase. In Fig. 4, we show the impact of dynamical electron-phonon scattering on the electron and phonon properties of 1T-TiSe2 when the system is just above the structural instability . Temperatures used in the theoretical simulations are scaled with respect to the value from density functional perturbation theory (DFPT), which is 1105 K. The DFPT is the electronic temperature at which the acoustic mode at the M point becomes soft in the harmonic Born-Oppenheimer approximation. Panels (a)-(d) show the electron spectral functions along the -M high symmetry direction and CECs (at Fermi energy) in the first Brillouin zone. The results for a temperature above but close to the CDW instability are depicted in Fig. 4 (a) and (b), while the other two panels are for , when the CDW fluctuating replica are gone and system enters into the normal phase. The electron-phonon interaction is included into the spectral function by using the dynamical Fan-Migdal electron self-energy Giustino (2017); Lee et al. (2023) based on inputs derived within the density functional perturbation theory Baroni et al. (2001); Giannozzi et al. (2017). Even though we use the original unit cell with no CDW periodic lattice distortions, we observe strong renormalizations of electron bands for very similar to the characteristic modifications observed in the CDW state, i.e., (i) finite spectral weight below the Ti 3d conduction band at the M point, coming from replicas of Se 4p states at the point, (ii) the flower-shaped contour at the point close to the Fermi level coming from the replicas of the three non-equivalent Ti 3d M pockets, and (iii) the loss of spectral weight at the particular momentum of the M pocket close to the Fermi level [see Fig. 1 (c)]. These signatures come from the electron-phonon fluctuations and very much resemble the CDW fluctuations observed in our ARPES data at RT. With the rise of the temperature, these electron-phonon fluctuations are reduced [see Fig. 4(e) and Supplementary Figure 2 in SI], and the electronic band structure becomes closer to the bare DFT result [red curves in Fig. 4(c)]. Note that the bimodal intensity profile in our simulations disappears continuously, while it is completely vanished above =1600 K, which is 500 K above the theoretical . We observe that these features are related to the soft acoustic phonon mode at the M point [see Fig.4(f)] and are strongest when the energy of this mode is close to zero, while the electron-phonon coupling is the strongest (see Supplementary Figure 2 in SI). When the temperature is elevated, the M-point acoustic phonon hardens, which quenches the electron-phonon scatterings between Se 4p and Ti 3d states. This further supports the correlation discussed above between the melting and recovery dynamics of the CDW fluctuating Se 4p band and hardening of the M phonon mode Otto et al. (2021). Similar microscopic explanation, which includes an active soft bosonic mode, was used many years ago to explain the fluctuating (or so-called pseudogap) region above the CDW phase Sadovskii (2001) in the quasi-1D Peierls systems Lee et al. (1973); McKenzie (1995), as well as in the high- cuprates Pines (1997), and more recently in metallic 2H-NbSe2 and 2H-TaSe2 Kuchinskii et al. (2012). Further, spectral weight redistribution around the Fermi level due to electron-phonon-related fluctuations for was once discussed for 1T-TiSe2 Yoshiyama et al. (1986), but it was never thoroughly appreciated and later connected to the replicas and other photoemission signatures in 1T-TiSe2. Our microscopic interpretation is also in line with the observation that the CDW fluctuations reach a maximum at , as observed with X-ray thermal diffuse scattering Holt et al. (2001). The temperature dependence of the phonon scattering lifetime characterized by hardening of the soft phonon mode with increasing temperature and a concomitant reduction of the electron scattering rates (i.e., ), consistent with transport measurements Velebit et al. (2016), is shown in Supplementary Figures 2 and 3. It shows the proportionality between the larger CDW gap and the longer lifetime of the amplitude mode. This is most likely caused by less quasiparticle damping and suggests that even in the short-range ordered CDW regions, there is a finite CDW gap that suppresses the damping of the coherent (and re-hardened) amplitude mode oscillation. Notably, the figures also show that artificial introduction of small periodic lattice distortions leads to an increase in the amplitude mode lifetime and better lifetime agreement with the experiment. However, the dynamical electron-phonon scattering alone redistributes the density of states enough to account for a pseudogap formation, as shown in Supplementary Figure 3 in SI. We therefore conclude that while static CDW gaps vanish above the transition temperature, the dynamic pseudogap persists. This affects both the electronic scattering and phonon suppression rates.
Figure 4(f) additionally shows the phonon spectral function of 1T-TiSe2 in the fluctuation regime with included phonon-electron coupling, where part of the M phonon spectral weight extends up to negative frequencies. This indicates that in the fluctuation region where , the atomic motions fluctuate between normal and CDW states. Note that equilibrium electron diffraction measurements Cheng et al. (2022) as well as X-ray diffuse scattering Woo et al. (1976b); Guo et al. (2025) observed both L and M diffusive peaks at RT, suggesting the presence of the fluctuating periodic lattice distortions above . In Supplementary Figure 4 we demonstrate how the dynamical electron-phonon scattering introduces very similar modification to the electronic structure as the symmetry-breaking periodic lattice distortions in the CDW state. The system could, therefore, support the excitation of the CDW amplitude mode at , as it is observed experimentally following optical photoexcitation [Fig. 3(f)]. These theoretical results demonstrate that the features (i)-(iii) as well as the CDW amplitude mode observed in the CDW fluctuation regime could be explained by the dynamical electron-phonon effects (also supported by our scanning tunneling microscopy measurements - see Supplementary Figure 5 in SI), which opposes the common belief that the fluctuation regime is dominated by the electron-hole correlations Monney et al. (2012, 2015). The latter mechanism could still be present in the formation process of the CDW fluctuations in 1T-TiSe2, but more likely as a higher-order correcting effect. This is further corroborated by the fact that the equivalent electron-related collective soft mode (e.g., plasmon or exciton), which could produce the above features in scattering with Ti 3d or Se 4p states, is not observed in the normal phase above , neither in experiments Lin et al. (2022); Kogar et al. (2017) nor theoretical studies Lian et al. (2019).
In conclusion, our study provides direct evidence of charge density wave (CDW) fluctuations in 1T-TiSe2 at room temperature using time-resolved extreme ultraviolet (XUV) momentum microscopy. We observe the persistence of the backfolded Se 4p band well above , demonstrating that CDW fluctuations remain detectable even at elevated temperatures. The ultrafast melting and recovery of these fluctuations, featuring the presence of coherent CDW amplitude phonon modes, highlight the importance of electron-phonon interactions in stabilizing the CDW phase at elevated temperatures.
Our experimental results, combined with density functional theory (DFT) calculations incorporating electron self-energy corrections, show that the electron-phonon interaction remains the dominant contributor to the CDW mechanism at room temperature. The observation of coherent amplitude phonon oscillations further supports this assertion, challenging the notion that the CDW fluctuating phase is purely driven by excitonic correlations. Instead, our findings suggest that while electron-hole coupling could also play a role, electron-phonon interactions are essential for the persistence of CDW fluctuations above with strong implications for the low-temperature CDW phase formation.
These results not only refine the understanding of CDW physics in 1T-TiSe2 but also have broader consequences for the study of fluctuating phases in other transition metal dichalcogenides and strongly correlated materials. Given the emerging connection between CDW fluctuations and unconventional superconductivity, our findings may provide valuable insights into the interplay between these two phenomena. Future studies leveraging advanced ultrafast techniques could further elucidate the microscopic mechanisms governing the fluctuating CDW phase, potentially unveiling new pathways for manipulating correlated electron systems in quantum materials.
Methods
Experimental Methods. The experimental apparatus is centered around a custom-designed, polarization-tunable ultrafast XUV beamline Comby et al. (2022), interfaced with a Momentum Microscope end-station Medjanik et al. (2017). This beamline is driven by a commercial high-repetition-rate Yb fiber laser (1030 nm, 135 fs FWHM, 50 W) from Amplitude Laser Group, operating at 166 kHz. In the probe branch, a segment of the laser output is frequency-doubled using a BBO crystal, yielding 515 nm pulses (5 W) that are tightly focused into a thin, high-density argon gas jet configured in an annular geometry to facilitate high-order harmonic generation (HHG). Spatial filtering through pinholes effectively separates the intense 515 nm fundamental from the emerging XUV beamlet. To extract the 9th harmonic (21.6 eV), a combination of SiC/Mg multilayer mirrors (NTTAT) and a 200 nm Sn metallic filter is used for spectral selection. On the pump side, a small portion of the original 1030 nm laser pulse (135 fs FWHM) is directed as the IR excitation. The IR pump and XUV probe pulses are recombined collinearly via a drilled mirror and focused onto the target. Bulk iodine vapor transport-grown - single crystals Watson et al. (2019) are freshly cleaved and positioned on a motorized 6-axis hexapod for high-precision alignment within the main vacuum chamber, which maintains an ultra-high base pressure of 210-10 mbar. Photoemission data acquisition is carried out with a custom-designed time-of-flight momentum microscope featuring an advanced front lens capable of multiple operating modes (GST mbH) Tkach et al. (2024b). This detection system enables simultaneous mapping of the entire surface Brillouin zone over a broad binding energy range without necessitating sample orientation adjustments Medjanik et al. (2017). Measurements in this study were conducted at room temperature, where the total energy resolution reaches . This resolution, , results from the convolution of the Fermi-Dirac distribution width at a given temperature , the bandwidth of the ionizing radiation , and the time-of-flight detector resolution . The thermal broadening of the Fermi-Dirac distribution is estimated as , corresponding to approximately at room temperature. For data post-processing, an open-source workflow Xian et al. (2020, 2019) is employed to efficiently convert raw single-event-based datasets into calibrated, binned data hypervolumes, incorporating appropriate axis calibration. Further details about the experimental setup are provided in Ref. Fragkos et al. (2025).
For STM, the surface was cleaved in situ and directly measured at room temperature in ultrahigh vacuum with a base pressure below mbar. STM measurements were performed with an Omicron LT-STM instrument and performed in constant current mode by applying a bias voltage to the sample.
The native defects observed with STM on the sample perfectly match those observed in Arpaia et al. (2019), with an estimated defect density of . Only very few defects of type D associated to intercalated Ti atoms could be observed (density ), validating the low Ti doping of the sample. No signs of CDW modulation have been observed with STM even close to [Supplementary Figures 5(a) and 5(b) in SI]. Indeed, the FFT-amplitude plot [Supplementary Figure 5(c)] only shows spots corresponding to the unit cell charge modulation.
Computational Methods. Ground state calculations were performed within the framework of the DFT using the Quantum Espresso package Giannozzi et al. (2017). We use optimized norm-conserving Vanderbilt pseudopotentials Hamann (2013) with the PBE exchange-correlation functional Perdew et al. (1996) from the PseudoDojo library van Setten et al. (2018). The plane wave energy cutoff was fixed to Ry. To reduce the interaction between the periodic images we modeled the 2D layers in periodic cells with Å vacuum in the direction normal to the layer. The convergence criterion for energy is set to 10-8 eV and the atomic positions are relaxed until the Hellmann-Feynman forces are less than 10-4 eV/Å. A lattice parameter of is obtained with the PBE functional. The relaxed atomic positions in fractional (crystal) coordinates are Ti (0.0000000000, 0.0000000000, 0.5000000000), Se (0.3333333333, 0.6666666667, 0.4163666236), and Se (0.6666666667, 0.3333333333, 0.5836333764). To obtain the ground state charge density the 2D Brillouin zone is sampled with k-point mesh. In order to determine the , i.e., the temperature at which the relevant acoustic phonon mode becomes unstable, we use the Fermi-Dirac smearing functions with the electronic temperatures ranging from 100 K to 1600 K, while for the calculations of the electron self-energy we range from 1100 K to 1600 K.
For the electron-phonon-coupling calculation, we use maximally localized Wannier functions Marzari et al. (2012), with 11 initial projections of d orbitals on the Ti sites and p orbitals on the Se atom sites, which have been implemented in the EPW code Lee et al. (2023). The electronic structure, dynamical matrices, and electron-phonon matrix elements were obtained from DFT and density functional perturbation theory (DFPT) Baroni et al. (2001) calculations, and they were used as the initial data for electron-phonon coupling. Initial DFPT phonon calculations were done on corse grids of and for the undistorted original unit cell, as well as for various mentioned above. The soft acoustic phonon mode goes to zero at K. In order to get a more accurate transition temperature one would need to include the anharmonic corrections Zhou et al. (2020).
To account for the dynamical renormalization of the electronic structure due to electron-phonon interactions we calculate the Fan-Migdal electron self-energy Giustino (2017)
where are the electron-phonon matrix elements, are the DFT electron energies, are the phonon frequencies, while and are the Bose-Einstein and Fermi-Dirac distribution functions. The fine momentum grid of was used for the summation in Eq. (1), while was set to 30 meV. The electron spectral function is calculated as
In the CDW systems, the input parameters entering Eqs. (1) and (2), such as and depend on the temperature (e.g., the soft M phonon mode softens as one approaches ), therefore, is calculated for several temperatures close and far from .
The dynamical phonon spectral function is calculated as
where is the phonon self-energy accounting for the dynamical renormalizations of the phonons due to coupling with electrons. The real part of the phonon self-energy shifts the DFPT frequencies , while the imaginary part introduces the broadening . The phonon self-energy due to electron-phonon coupling is calculated as
The summation of the latter equation was done on momentum grid.
Data Availability
The ARPES and DFT data that support the findings of this study are available in Zenodo under the identifier https://doi.org/10.5281/zenodo.17813602. Additional data can be obtained from the authors upon reasonable request.
Acknowledgements.
We thank Claude Monney for useful discussions. We thank Nikita Fedorov, Romain Delos, Pierre Hericourt, Rodrigue Bouillaud, Laurent Merzeau, and Frank Blais for technical assistance and Aymen Mahmoudi, Joel Morf, and Frederic Chassot for their assistance during STM measurements. J.S., S.S., and J.M. would like to thank the QM4ST project with Reg. No. CZ.02.01.01/00/22_008/0004572, cofunded by the ERDF as part of the MŠMT. The research leading to these results has received funding from LASERLAB-EUROPE (grant agreement no. 871124, European Union’s Horizon 2020 research and innovation programme). We acknowledge the financial support of the IdEx University of Bordeaux / Grand Research Program "GPR LIGHT". We acknowledge support from ERC Starting Grant ERC-2022-STG No.101076639, Quantum Matter Bordeaux, AAP CNRS Tremplin, and AAP SMR from Université de Bordeaux. This work is part of the ULTRAFAST and TORNADO projects of PEPR LUMA and was supported by the French National Research Agency, as a part of the France 2030 program, under grants ANR-23-EXLU-0002 and ANR-23-EXLU-0004. SF acknowledges funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie 2024 Postdoctoral Fellowship No 101198277 (TopQMat). JS acknowledges funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska‑Curie grant agreement No 101209345—ART.QM funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work was funded by the Würzburg-Dresden Cluster of Excellence on Complexity, Topology and Dynamics in Quantum Matter – ctd.qmat (EXC 2147, project-id 390858490) and by the DFG through SFB1170 “Tocotronics” and directly via RE 1469/13-2. N.G.E., and D.N. acknowledge financial support from the project "Podizanje znanstvene izvrsnosti Centra za napredne laserske tehnike (CALTboost)" financed by the European Union through the National Recovery and Resilience Plan 2021-2026 (NRPP), European Regional Development Fund for the project ‘Materials for clean energy, advanced sensors and quantum technologies’ (Grant No. PK.1.1.10.0002), and Croatian Science Foundation (Grant no. IP-2025-02-5926). QC also acknowledges support from the TERAQUANTUM project of the Région Nouvelle-Aquitaine. This work was supported by the TWISTnSHINE , funded as project No. LL2314 by Programme ERC CZ.Author contributions
S.F., H.O., S.B., and J.S. performed the experiments, supported by S.S., A.J., Q.C., Y.M., and M.P.T.M.. S.F., H.O., S.B., and J.S. analyzed the experimental data. D.D. and S.P. participated in maintaining the laser system. N.G.E. and D.N. performed the theoretical calculations. F.D. and K.R. synthesized and characterized the samples. B.H. and J.S. performed the STM measurements. S.F., H.O., N.G.E., A.J., S.S., Q.C., M.P.T.M., M.Ü., F.D., B.H., D.D., S.P., F.B., J.M., Y.M., F.R., K.R., D.N., S.B., J.S. contributed to the interpretation and discussion of the results. J.S., S.B., D.N., H.O., and S.F. wrote the manuscript with inputs from coauthors. J.S., S.B., and D.N. conceived and planned the project.
Competing interests
The authors declare no competing interests.